Exact solution

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 1)

First order Upwind

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 2)

Lax Wendroff, 25 time steps

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 3)

Lax Wendroff

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 4)

Amplification Upwind

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 5)

Amplification Lax Wendroff

Examples Lecture 14; Jos van Kan, DIAM, Delft, December 2005 (slide 6)